1) Construct plots that show the wavelength-dependent energy spectrum of a blackbody at a temperature of 5800 K (approx. temperature of the Sun) using both the Planck distribution and the Raleigh-Jeans distribution. Confirm agreement between the two at long wavelength. a. What is the maximum emission wavelength at this temperature? b. What is the total power output 2) c. Using the Planck distribution, estimate what fraction of the Suns total power output is emitted in visible wavelengths (400-750 nm)
Answer #1

(a)

A should be weawned in nK 2.3 へ2/ 녀 Planka laud h 664x 13 -34 2x 10 2S x 10 2 0 LC

The plot is

Energy Density v/s Wavelength 12 Rayleigh Plancks_ 10 5000 0000 15000 20000 25000 30000 Wavelength(nm)

Gnuplot is used for plotting

The text document is

set zeroaxis
set xlabel 'Wavelength(nm)'
set ylabel 'energy density(J/kg m^3)'
set title 'Energy Density v/s Wavelength'
set term png
set output 'rp.png'
r(x)=(8*pi*8*10**(15))/x**4
p(x)=((8*pi*2*10**(20))/x**5)*(1/(exp(2.5*10**(4)/x)-1))
plot [1000:32000][0.1:12] r(x) title 'Rayleigh',p(x) title 'Plancks'


Write all this in a txt file and in gnuplot write

load 'filename.txt'

From graph one can see:

At higher wavelength i.e. (2500 nm- 3000 nm) there is agreement between Plancks and Rayleigh

maximum wavelength at this tempertaure is

\lambda_{max}=5000\text{nm}

(b) Total power emitted

Q=\sigma T^4

where \sigma is stefans constant

Q=5.6\times 10^{-8} \times 5800^4

Q=63.37\times 10^6\mathrm{J/m^2 s}